# Generalized $K$-theoretic invariants and wall-crossing via non-abelian localization

### v1

@@ 1.1.4. Comparison with Joyce–Liu invariants
We expect canonical homological lifts of $\varepsilon$ classes to exist in general, without assuming the existence of a framing functor, and we hope to address this question in the future. In Section 8.4 we explain that our wall-crossing formulas for $\varepsilon$ invariants implies Joyce's wall-crossing formulas for the homological lifts, at least if we consider only intersection numbers of (algebraic) tautological classes.
@@ 1.2. Analogy with motivic wall-crossing
A remarkable theorem of Joyce [Joy4, Theorem 8.7] is that the $\varepsilon$ invariants no longer have a pole at $q=1$ – see also [BNK] and [BR] for more recent approaches to the proof of this result. More precisely, they are regular elements, i.e. lie in
@@ 3.1. $K$-theory of stacks
Remark 3.3.
Instead of working with algebraic $K$-theory, another option is to use topological $K$-theory, in the sense of Blanc [Bla], and define
$K^{\ast}_{\textup{top}}({\mathfrak{X}})=K^{\textup{top}}_{0}(\mathsf{Perf}({\mathfrak{X}}))\oplus K^{\textup{top}}_{1}(\mathsf{Perf}({\mathfrak{X}}))\,.$
We can do the same with $G$ theory, and these have the same functoriality properties as their algebraic counterparts. It seems to us that everything in this paper could be rewritten in that language. Halpern-Leistner has kindly informed us that his non-abelian localization theorem [HL5] also holds for Blanc's $K$-theory. The advantage of doing so is that it makes the comparison with cohomology easier (cf. Section 8, and in particular Remark 8.14) since, for example, the Chern character induces an isomorphism between Blanc's $K$-theory and cohomology (with $\mathbb{Q}$}-coefficients) for separated finite type schemes.
@@ 3.1.1. Operations in $K$ and $G$ theory
There are some situations where pushforwards exist even without the map being proper, see [HLP]. A particularly important one is the following:
@@ 3.2. $K$-homology
There are two more conditions in the definition of $K$-homology in [Liu]. The ``equivariant localization'' is irrelevant for us since we are not working equivariantly. The ``finiteness condition'', on the other, is a fundamental difference and we really need to exclude it. For example, for us
@@ 5.2. The non-abelian virtual localization theorem
be the set of possible $\mu$-HN types. Any object in $\mathcal{A}$} has an associated $\mu$-HN type $c=(\alpha_{1},\ldots,\alpha_{n})$ where $\alpha_{i}=[E_{i}/E_{i-1}]$ are the types of its $\mu$-HN factors in (2). Given $c\in\mathrm{HN}_{\alpha}(\mu)$, we consider the substack ${\mathcal{M}}_{=c}^{\mu}\subseteq{\mathcal{M}}_{\alpha}$ of objects whose $\mu$-HN type is $c$.
@@
We say that there is a pseudo $\Theta$-stratification of ${\mathcal{M}}_{\alpha}$ adapted to $\mu$ if there is a partial ordering of $\mathrm{HN}_{\alpha}(\mu)$ which satisfies [^9] In the right hand side of (16), $\alpha_{i}+\alpha_{j}$ is put in a position that makes it a $\mu$-HN type, i.e. it is put between $\alpha_{k}$ and $\alpha_{k+1}$ such that $\mu(\alpha_{k})>\mu(\alpha_{i}+\alpha_{j})>\mu(\alpha_{k+1})$. In the case that $\mu(\alpha_{k})=\mu(\alpha_{i}+\alpha_{j})$ for some $k$, the right hand side should be interpreted as $(\ldots,\alpha_{i}+\alpha_{k}+\alpha_{j},\ldots)$.
@@
and for which
${\mathcal{M}}_{\leq c}^{\mu}\coloneqq\bigcup_{c^{\prime}\leq c}{\mathcal{M}}_{=c^{\prime}}^{\mu}$
forms a pseudo $\Theta$-stratification.
@@ 5.3.2. Derived enhancement
We explain now how to upgrade the statements above to the derived setting. The stacks of filtered and graded points of a derived stack inherit a natural derived enhancement (cf. [HLP, Theorem 5.1.1] or [HL3, Theorem 1.2.1]).
@@ 8.4. Cohomological descendents from $K$-theory descendents
Remark 8.14.
If one replaces algebraic $K$-theory by Blanc's topological $K$-theory (cf. Remark 3.3) then the same argument works for any $D\in H^{\ast}_{\textup{taut}}({\mathfrak{M}})$ since the Chern character defines an isomorphism between $K^{\ast}_{\textup{top}}(X^{n})$ and $H^{\ast}(X^{n})$. We expect that equality in Blanc's $K$-theory actually implies $A=B$.

### v2

@@ 1.1.4. Comparison with Joyce–Liu invariants
We expect canonical homological lifts of $\varepsilon$ classes to exist in general, without assuming the existence of a framing functor, and we hope to address this question in the future. In Section 8.4 we explain that our wall-crossing formulas for $\varepsilon$ invariants implies Joyce's wall-crossing formulas for the homological lifts, at least if we consider only intersection numbers of (algebraic) tautological classes. In Appendix A we discuss a version of the entire theory where we replace algebraic $K$-theory by other additive invariants of dg categories, such as Hochschild homology or Blanc's topological $K$-theory. Topological $K$-theory, in particular, is closer to cohomology and allows us to deduce wall-crossing formulas without algebraicity assumptions.
@@ 1.2. Analogy with motivic wall-crossing
A remarkable theorem of Joyce [Joy4, Theorem 8.7] is that the $\varepsilon$ invariants no longer have a pole at $q=1$ – see also [BHLNK] and [BR] for more recent approaches to the proof of this result. More precisely, they are regular elements, i.e. lie in
@@ 1.3. Conjectures and future directions
Let us finally mention the following curious question (as usual, due to Dominic Joyce and his collaborators): since, as we explain, the present paper develops precisely the machinery required for the wall-crossing in the derived categories, the Remark 4.3.2 (c) in [GJT] hereby starts making sense; in a future, we would like to investigate whether, as loc. cit. predicts, one can, thus, introduce ``Joyce structures with descendants'', say, for projective surfaces.
@@ 3.1.1. Operations in $K$ and $G$ theory
There are some situations where pushforwards exist even without the map being proper, see [HLP2]. A particularly important one is the following:
@@ 3.2. $K$-homology
There are two more conditions in the definition of $K$-homology in [Liu]. The ``equivariant localization'' is irrelevant for us since we are not working equivariantly. The ``finiteness condition'', on the other hand, is a fundamental difference and we really need to exclude it. For example, for us
@@ 5.2. The non-abelian virtual localization theorem
be the set of possible $\mu$-HN types. Any object in $\mathcal{A}$} has an associated $\mu$-HN type $c=(\alpha_{1},\ldots,\alpha_{n})$ where $\alpha_{i}=[E_{i}/E_{i-1}]$ are the types of its $\mu$-HN factors in (2).
@@
We say that there is a pseudo $\Theta$-stratification of ${\mathcal{M}}_{\alpha}$ adapted to $\mu$ if there is a partial ordering $\geq$ of $\mathrm{HN}_{\alpha}(\mu)$ which satisfies [^9] In the right hand side of (16), $\alpha_{i}+\alpha_{j}$ is put in a position that makes it a $\mu$-HN type, i.e. it is put between $\alpha_{k}$ and $\alpha_{k+1}$ such that $\mu(\alpha_{k})>\mu(\alpha_{i}+\alpha_{j})>\mu(\alpha_{k+1})$. In the case that $\mu(\alpha_{k})=\mu(\alpha_{i}+\alpha_{j})$ for some $k$, the right hand side should be interpreted as $(\ldots,\alpha_{i}+\alpha_{k}+\alpha_{j},\ldots)$.
@@
and a pseudo $\Theta$-stratification of ${\mathcal{M}}_{\alpha}$ indexed by $(\mathrm{HN}_{\alpha}(\mu),\geq)$ for which the $\mathbb{C}$}-points of ${\mathcal{M}}_{=c}^{\mu}$ correspond to objects of $\mathcal{A}$} of $\mu$-HN type $c$.
@@ 5.3.2. Derived enhancement
We explain now how to upgrade the statements above to the derived setting. The stacks of filtered and graded points of a derived stack inherit a natural derived enhancement (cf. [HLP2, Theorem 5.1.1] or [HL3, Theorem 1.2.1]).
@@ 8.4. Cohomological descendents from $K$-theory descendents
We will see in Appendix A that a better version of the previous proposition can be achieved if we replace algebraic $K$-theory by topological $K$-theory.
Appendix A $\mathsf{E}$-Hall algebra for additive invariant $\mathsf{E}$
In this section we will explain how the material in the present paper generalizes if we replace algebraic $K$-theory by other additive invariants, such as Hochschild homology or Blanc's topological $K$-theory.
A.1. Additive invariants
Let $R$ be a ring and let $\mathsf{Mod}_{R}$ be the category of $R$-modules. An additive invariant with values in $\mathsf{Mod}_{R}$, in the sense of Tabuada [Tab], is a functor $\mathsf{E}\colon\mathsf{dgcat}\to\mathsf{Mod}_{R}$ from the category of ($\mathbb{C}$}-linear) dg categories to the category of $R$-modules which satisfies 2 properties: the image by $\mathsf{E}$ of a Morita equivalence of dg categories is an isomorphism, and if $\mathcal{C}=\langle\mathcal{C}_{1},\mathcal{C}_{2}\rangle$ is a semi-orthogonal decomposition then $\mathsf{E}({\mathcal{C}}_{1})\oplus\mathsf{E}({\mathcal{C}}_{2})\xrightarrow{\sim}\mathsf{E}({\mathcal{C}})$. The most basic example of an additive invariant is $K_{0}$ (with $R={\mathbb{Z}}$).
By [Tab, Théorème 6.1.], $\mathsf{E}$ being additive is equivalent to factoring through
$\mathsf{dgcat}\to\mathsf{Hmo}_{0}\to\mathsf{Mod}_{R}$
where $\mathsf{Hmo}_{0}$ is a category whose objects are the same as $\mathsf{dgcat}$ and
$\mathrm{Hom}_{\mathsf{Hmo}_{0}}({\mathcal{C}},{\mathcal{D}})=K_{0}\big(\mathrm{rep}({\mathcal{C}},{\mathcal{D}})\big)\,.$
This means that if $F\colon{\mathcal{C}}\to{\mathcal{D}}$ is a dg functor then $\mathsf{E}(F)$ is the image of $F]\in K_{0}\big(\mathrm{Fun}^{\textup{ex}}({\mathcal{C}},{\mathcal{D}})\big)$ through the composition
$K_{0}\big(\mathrm{Fun}^{\textup{ex}}({\mathcal{C}},{\mathcal{D}})\big)\to K_{0}\big(\mathrm{rep}({\mathcal{C}},{\mathcal{D}})\big)\to\mathrm{Hom}_{R}(\mathsf{E}({\mathcal{C}}),\mathsf{E}({\mathcal{D}}))\,$,
and in particular if $0\to F_{1}\to F_{2}\to F_{3}\to 0$ is an exact sequence of dg functors ${\mathcal{C}}\to{\mathcal{D}}$ then $\mathsf{E}(F_{2})=\mathsf{E}(F_{1})+\mathsf{E}(F_{3})$.
Any additive invariant is a module over $K_{0}$, in the sense that for any two dg categories $\mathcal{C}$} and $\mathcal{D}$} we have a canonical morphism of $R$-modules
(37) $K_{0}(\mathcal{C})\otimes_{\mathbb{Z}}\mathsf{E}(\mathcal{D})\to\mathsf{E}(\mathcal{C}\otimes\mathcal{D})\,.$
Indeed, this morphism is obtained by adjunction from
$K_{0}({\mathcal{C}})=K_{0}(\textup{Fun}^{\textup{ex}}({\bf 1},{\mathcal{C}}))\xrightarrow{\otimes\operatorname{id}_{{\mathcal{D}}}}K_{0}(\textup{Fun}^{\textup{ex}}({\mathcal{D}},{\mathcal{C}}\otimes{\mathcal{D}}))\to\mathrm{Hom}_{R}(\mathsf{E}({\mathcal{D}}),\mathsf{E}({\mathcal{C}}\otimes{\mathcal{D}}))\,.$
We say that $\mathsf{E}$ is lax monoidal if it comes together with $R$-module homomorphisms
$R$ $\to\mathsf{E}(\mathsf{Perf}({\mathsf{pt}}))$
$\mathsf{E}({\mathcal{C}})\otimes_{R}\mathsf{E}({\mathcal{D}})$ $\to E({\mathcal{C}}\otimes{\mathcal{D}})$
satisfying some natural compatibilities, and we say that $\mathsf{E}$ is (strongly) monoidal if these homomorphisms are isomorphisms.
Example A.1.
Hochschild homology $HH(\mathcal{{\mathcal{C}}})=\bigoplus_{i\in{\mathbb{Z}}}HH_{i}(\mathcal{C})$ is a strongly monoidal additive invariant with values in $\mathbb{C}$} vector spaces. Other variations such as cyclic homology or periodic cyclic homology are also monoidal additive invariants. On the other hand, $HH_{0}(-)$ is lax monoidal but not strongly monoidal.
A.2. $\mathsf{E}$-Hall algebra and $\varepsilon$ invariants
Let $\mathsf{E}$ be an additive invariant which we will further assume that commutes with infinite direct sums (this is required for the analog of Proposition 3.7 to hold for $\mathsf{E}$).
We define for a derived stack $\mathfrak{X}$}
$\mathsf{E}^{\ast}({\mathfrak{X}})=\mathsf{E}(\mathsf{Perf}({\mathfrak{X}}))\,.$
Just as $K^{\ast}$, $\mathsf{E}^{\ast}$ admits arbitrary pullbacks and pushforwards along maps of stacks whose pushforward preserves perfect complexes. There is a tensor product map
[…]
When $\mathsf{E}$ is lax monoidal there is also a tensor product map $\mathsf{E}^{\ast}({\mathfrak{X}})\otimes_{R}\mathsf{E}^{\ast}({\mathfrak{X}})\to\mathsf{E}^{\ast}({\mathfrak{X}})$.
We may define the $\mathsf{E}$-homology group $\mathsf{E}_{\ast}({\mathfrak{X}})$ in complete analogy with Definition 3.8, with the caveat that there are two possible slightly different variations: if $\mathsf{E}$ is a general additive invariant, an element of $\mathsf{E}_{\ast}({\mathfrak{X}})$ is defined to be a collection of functionals
$\{\phi_{S}\colon\mathsf{E}^{\ast}({\mathfrak{X}}\times S)\to\mathsf{E}^{\ast}(S)\}$
which are $K^{\ast}(S)\otimes_{\mathbb{Z}}R$-linear, with the same compatibility requirement. If $\mathsf{E}$ is lax monoidal it seems more natural to require the functionals to be $\mathsf{E}^{\ast}(S)$-linear, or otherwise $\mathsf{E}_{\ast}({\mathfrak{X}})$ becomes potentially very large. Either option provides a reasonable theory, so in what follows we will leave ambiguous which option we take. If $\mathsf{E}$ is strongly monoidal and we require $\mathsf{E}^{\ast}(S)$-linearity then
$\mathsf{E}_{\ast}({\mathfrak{X}})\simeq\mathrm{Hom}_{R}(\mathsf{E}^{\ast}({\mathfrak{X}}),R)\,$
is just a space of functionals.
We have analogs of all the functoriality properties of Section 3.2. For example, for any additive invariant we have a cap product
$\cap\colon K^{\ast}({\mathfrak{X}})\otimes\mathsf{E}_{\ast}({\mathfrak{X}})\to\mathsf{E}_{\ast}({\mathfrak{X}})$
obtained by dualizing (38). The analog of (6) is still a map $G({\mathfrak{X}})\to\mathsf{E}_{\ast}({\mathfrak{X}})$. Indeed, (6) is obtained from Lemma 3.10, which gives a functor $D^{b}\mathrm{Coh}({\mathfrak{X}})\otimes\mathsf{Perf}({\mathfrak{X}}\times S)\to\mathsf{Perf}(S)$, and we can use it together with (37) to obtain
$G({\mathfrak{X}})\otimes\mathsf{E}^{\ast}({\mathfrak{X}}\times S)=K_{0}(D^{b}\mathrm{Coh}({\mathfrak{X}}))\otimes\mathsf{E}(\mathsf{Perf}({\mathfrak{X}}\times S))$ $\to\mathsf{E}(D^{b}\mathrm{Coh}({\mathfrak{X}})\otimes\mathsf{Perf}({\mathfrak{X}}\times S))$
$\to\mathsf{E}(\mathsf{Perf}(S))=\mathsf{E}^{\ast}(S)\,.$
In particular, we get fundamental classes ${\mathfrak{X}}]_{\mathsf{E}}\in\mathsf{E}_{\ast}({\mathfrak{X}})$ for stacks as in Theorem 3.11.
With this in place, all the material from Section 4 is adapted to additive invariants $\mathsf{E}$ in a straightforward way, and we obtain a $\mathsf{E}$-Hall algebra
$\mathbb{E}({\mathcal{A}})\coloneqq\mathsf{E}_{\ast}({\mathfrak{M}}_{{\mathcal{A}}}^{\mathrm{rig}})$
and invariants
$\delta_{\alpha}^{\mu,\mathsf{E}},\varepsilon_{\alpha}^{\mu,\mathsf{E}}\in\mathbb{E}({\mathcal{A}})\,.$
For $\varepsilon$ invariants we need to either assume that $R$ contains $\mathbb{Q}$} or extend to a ring containing $\mathbb{Q}$} as in Definition 4.2.
A.3. Non-abelian localization
The last ingredient in our setup is the non-abelian localization formula for additive invariants, which will soon appear in a revision of [HL5]. Once we have it, it follows by the same reasoning that the invariants $\delta^{\mu,\mathsf{E}}_{\alpha},\varepsilon^{\mu,\mathsf{E}}_{\alpha}$ satisfy the exact same wall-crossing formulas (20) and (21). We explain the statement here.
To imitate the notation used in $K$-theory we denote by
$\chi_{\mathsf{E}}({\mathfrak{X}},-)\colon\mathsf{E}^{\ast}({\mathfrak{X}})\to\mathsf{E}^{\ast}({\mathsf{pt}})\,$
the map induced by the pushforward morphism $\mathsf{Perf}({\mathfrak{X}})\to\mathsf{Perf}({\mathsf{pt}})=D^{b}\mathrm{Coh}({\mathsf{pt}})$, where $\mathfrak{X}$} is a stack satisfying the hypothesis of Theorem 5.6.
Theorem A.2 ([HL5]).
Let $\mathfrak{X}$} be as in Theorem 5.6 and let $\mathsf{E}$ be an additive invariant. Then, for any $V\in\mathsf{E}^{\ast}({\mathfrak{X}})$, the following formula holds:
(39) $\chi_{\mathsf{E}}({\mathfrak{X}},V)=\chi_{\mathsf{E}}(\mathcal{{\mathfrak{X}}}^{\textup{ss}},V|_{\mathcal{{\mathfrak{X}}}^{ss}})+\sum_{c\in\Gamma\setminus\{0\}}\chi_{\mathsf{E}}({\mathfrak{Z}}_{c},V|_{{\mathfrak{Z}}_{c}}\otimes E_{c})\in\mathsf{E}^{\ast}({\mathsf{pt}})\,.$
A.4. Topological $K$-theory versus cohomology
In [Bla], Blanc defines topological $K$-theory of a dg category as a spectrum which, by taking homotopy groups, produces abelian groups $K^{\mathrm{top}}_{i}({\mathcal{C}})$ satisfying Bott periodicity $K_{i+2}^{\mathrm{top}}({\mathcal{C}})\simeq K_{i}^{\mathrm{top}}({\mathcal{C}})$. Then
$K^{\mathrm{top}}({\mathcal{C}})=K^{\mathrm{top}}_{0}({\mathcal{C}})\oplus K^{\mathrm{top}}_{1}({\mathcal{C}})$
is another example of a lax monoidal additive invariant (with $R={\mathbb{Z}}$). Let us denote by $K^{\ast}_{\textup{top}}({\mathfrak{X}}),K_{\ast}^{\textup{top}}({\mathfrak{X}})$ the topological $K$-theory of $\mathsf{Perf}({\mathfrak{X}})$ and topological $K$-homology of $\mathfrak{X}$}.
Topological $K$-theory admits a canonical map from algebraic $K$-theory $K_{0}(-)\to K_{0}^{\textup{top}}(-)$ and the Chern character map factors through
$\mathrm{ch}\colon K^{\ast}_{\textup{top}}({\mathfrak{X}})\to H^{\ast}({\mathfrak{X}})\,.$
If $X$ is a separated finite type scheme over $\mathbb{C}$} then the Chern character map is an isomorphism after tensoring with $\mathbb{Q}$} [Bla, Theorem 1.1.b]. Results in a similar direction for (global quotient) stacks can be found in [HLP1].
The entire content of Section 8 can be adapted to topological $K$-theory. In particular, we can talk about an element of $H_{\ast}({\mathfrak{M}}^{\mathrm{rig}}_{\mathcal{A}})$ being a homological lift of the $\varepsilon$ class
$\varepsilon_{\alpha}^{\mu,K_{\textup{top}}}\in K_{\ast}^{\textup{top}}({\mathfrak{M}}_{\mathcal{A}}^{\mathrm{rig}})\eqqcolon{\mathbb{K}}_{\textup{top}}({\mathcal{A}})\,.$
Since topological $K$-theory can ``see'' non-algebraic cohomology classes we get the following improvement of Proposition 8.13:
Proposition A.3.
Let $\mathfrak{M}$} be a moduli stack of complexes on a smooth projective variety $X$. Suppose that $A,B\in H_{\ast}({\mathfrak{M}})$ are homological lifts of the same class in $K_{\ast}^{\textup{top}}({\mathfrak{M}})$. Then $A$ and $B$ agree on tautological classes, i.e.
$\int_{A}D=\int_{B}D$
for any $D\in H^{\ast}_{\textup{taut}}({\mathfrak{M}})$.
Proof.
The proof goes exactly as in Proposition 8.13, but we further use the fact that $\mathrm{ch}\colon K^{\ast}_{\textup{top}}(X^{n})_{\mathbb{Q}}\to H^{\ast}(X^{n})$ is an isomorphism. ∎
We expect that having the same homological lift actually implies that the two classes in homology agree. Indeed, this would be a consequence of the following natural looking conjecture:
Conjecture A.4.
For any (reasonable) stack $\mathfrak{X}$} and any $N\geq 0$ the truncated Chern character map
$\mathrm{ch}_{\leq N}\colon K^{\ast}_{\textup{top}}({\mathfrak{X}})_{\mathbb{Q}}\to H^{\ast}({\mathfrak{X}})\twoheadrightarrow H^{\leq N}({\mathfrak{X}})$
is surjective.